- The building should be much cooler today because it has had time to catch up. However, there still could be a problematic areas until floors can be balanced starting next week. Until balancing, please try to keep blinds closed on the hot days to help with the heat. Should you have heat concerns today, please let us know.
DOED 3rd Floor East Tenant Move:
- Starts tomorrow night around 4:00 pm. Please ensure all items are packed and labeled for the move after hours.
Update of Construction From Last Night:
- moved glass top up to Thurman’s office on 6th floor,
- moved glass out of 764, pushed desk away from the wall in room 764.
- worked on grid in the east corridor of social security on 7,
- demo of the east walls for ODAR and put up plastic.
- tile drop on 8 east
- worked on drywall behind NE stair 3 on 9 and 8,
- framed drinking fountain insets and ceilings on 3,6,7
- Note** We inspected the ceilings in the restrooms on 2,3,6 and they are to drywall those tomorrow night. Tomorrow night they will stock drywall and install on 2nd floor.
- worked on 9th floor dampers in the men’s restroom,
- installed new merv 8 filters in the lobbies,
- Worked on slot diffusers on 6 east
- Installed the slot diffuser box on 8 north east.
- worked on controls on boxes in 3 north, programming boxes on 3rd floor,
- replacement of CO2 sensors on 6th floor
- worked on 6 east lights, 6 south lights, fire alarm rough in 7 east, furniture power put back on 6 south and 3 east.
- worked on data put backs on 6 south
- worked on 3rd floor east cubicle move back and 6 south.
- worked on final coats and touch ups on 6 south,
- 7 south cut in ceilings and walls on 6 east D.O.J. perimeter,
- second coats on 6 east north of D.O.J.,
- final painted last bit of 9 east just south of stair 3.
- worked on sprinkler heads on 6 east, 7 east, 2 north life safety upgrades,
- rim on east elevation,
- head install on 8 north east behind stair 3.
- worked on dusting and general cleaning throughout the floors.